OpenStudy (anonymous):

3 is a zero of f(x)=x^3-x^2-7x+3. Express f(x) as a product of linear factors.

OpenStudy (zzr0ck3r):

you need to divide the cubic by x-3

OpenStudy (zzr0ck3r):

i would google long division of polynomials its not so easy to give examples in latex, and I cant use the draw option on this computer..

OpenStudy (zzr0ck3r):

http://www.purplemath.com/modules/polydiv2.htm

OpenStudy (anonymous):

that comes out to f(x)=x^2-2x-1. Then what? That's where I get lost.

OpenStudy (zzr0ck3r):

now factor x^2-2x-1

OpenStudy (zzr0ck3r):

your answer will be like this (x-3)(x+a)(x+b) for some a and b

OpenStudy (zzr0ck3r):

hmm u sure thats right?

OpenStudy (anonymous):

Um I just saw that, one sec. Because that's not factorable

OpenStudy (anonymous):

f(x)=x^2+2x-1

OpenStudy (anonymous):

you have to factor using i or something?

OpenStudy (zzr0ck3r):

-(-x+sqrt(2)-1)(x+sqrt(2)+1)

OpenStudy (zzr0ck3r):

\(x^2+2x-1=0\iff (x+1)^2=1+1\iff x+1=\pm\sqrt{2}\iff x=\pm\sqrt{2}-1\)

OpenStudy (zzr0ck3r):

\(x^2+2x-1=0\iff (x+1)^2=1+1\iff x+1=\pm\sqrt{2}\iff x\\=\pm\sqrt{2}-1\)

OpenStudy (zzr0ck3r):

so we get \((x-(\sqrt{2}-1)(x-(-\sqrt{2}-1))\)

OpenStudy (zzr0ck3r):

I factored by completing the square...

OpenStudy (zzr0ck3r):

thus we get \((x-3)(x-\sqrt{2}+1)(x+\sqrt{2}+1)\)

OpenStudy (anonymous):

how do you get " -(-x+sqrt(2)-1)(x+sqrt(2)+1)"? I completed the square for f(x)=x^2+2x-1 and got (x+1)^2-2